The sample data below have been collected based on a simple random sample from a normally distributed population. Complete parts a and b 3 7 4 Compute a 98% confidence interval estimate for the population mean. The 98% confidence interval for the population mean is from 1 to (Round to two decimal places as needed. Use ascending order.)
Answer #1

Solution:

The mean and standard deviation for the given data is:

3 +3 +0 +3+2+7+9+4+8+4 10 4.3

s=sqrt{rac{sum (x-ar{x})^2}{n-1}}

(3-4.3)2 + (3-4.3)2 + (0-13)2 + + (4-4.3)2 10 1

=2.8304

Now, the 98% confidence interval for the population mean is:

ar{x} pm t_{rac{0.02}{2}}left ( rac{s}{sqrt{n}} ight )

Where:

toga = 2.821 is the t critical value at 0.02 significance level for df=n-1=10-1=9

Therefore, we have:

  4.3 ± 2.821 (2.8301 V10

4.3 ± 2.52

(4.3- 2.52, 4.3 + 2.52)

(1.78,6.82)

Therefore, the 98% confidence interval for the population mean is (1.78,6.82)